{Exercise 8.21 (Algorithmic)}

Consumption of alcoholic beverages by young women of drinking age has been increasing in the United Kingdom, the United States, and Europe (The Wall Street Journal, February 15, 2006). Data (annual consumption in liters) consistent with the findings reported in The Wall Street Journal article are shown for a sample of 20 European young women.

87 82 199 174 97
170 222 115 133 169
164 113 99 171 0
93 0 93 110 329

Assuming the population is roughly symmetric, calculate the sample mean and sample standard deviation. Then construct a 95% confidence interval for the mean annual consumption of alcoholic beverages by European young women (Round all answers to 1 decimal).

The mean is  and the standard deviation is

The confidence interval is (  ,  )
Answer #1

Solution

\small \mathbf{X} X-X (X-)
87 -44 1936
82 -49 2401
199 68 4624
174 43 1849
97 -34 1156
170 39 1521
222 91 8281
115 -16 256
133 2 4
169 38 1444
164 33 1089
113 -18 324
99 -32 1024
171 40 1600
0 -131 17161
93 -38 1444
0 -131 17161
93 -38 1444
110 -21 441
329 198 39204
X = 2620 Σ (x - X) = 104364

Mean (X) = ΣΧ2620 – 131.0 Γ 20.

Standard deviation (s) = 1 - (x – X)? _ ./104364 _ 0 -12 = V 20 -1 = 74.1

Degrees of freedom df) = n-1 = 20 -1 = 19

For 95% confidence level with degrees of freedom = 19

** = 2.093 (Refer t-table)

formula : Confidence interval for the population mean,

= X ****

=131.0 +2.093 * 74.1 ;

= 131.0±34.7

= (131.0 -34.7.131.0 +34.7)

= (96.3.165.7)

... The confidence interval is (96.3, 165.7)